You conducted an ocular inspection and you have noticed that a rectangular floor measuring 10m and 14m is needed to be made into rooms for patients.Make a model to illustrate the situation with appropriate mathematical solutions


1
Expert's answer
2021-03-17T10:27:22-0400

Each piece of brick pave is a square with an edge of 50 cm. How many pieces of brick paves will be needed to cover the rectangular floor that needs fixing?

"50\ cm=0.5\ m"


"\dfrac{10\ m}{0.5\ m}=20"

There are 20 pieces of brick paves in row.



"\dfrac{14\ m}{0.5\ m}=28"

There are 28 rows.


"20\times28=560"

560 pieces of brick paves will be needed to cover the rectangular floor.


Check

Area of one piece


"A_1=0.5\ m\times0.5\ m=0.25\ m^2"

Total area of the rectangular floor


"A_{floor}=10\ m\times14\ m=140\ m^2"

"\dfrac{140}{0.25}=560"

560 pieces of brick paves will be needed to cover the rectangular floor.